How many liters of 1% milk do you need to mix with 4% milk to make 20 liters of 2% milk
Answer by jorel1380(3719)

20x.02=.4
.01p+.04(20-p)=.4
p+4(20-p)=40
p+80-4p=40
80-3p=40
40=3p
p=40/3
You need 40/3 liters of 1% milk and 20/3 liters of 4% milk for your mix..